Two parallel metal plates with equal and opposite charges are separated by a distance x= 0.005 m. The electric potential at the positive plate is 23 V and that at the negative plate is 21 V. How large is the electric field anywhere between the plates?
1
Expert's answer
2020-04-23T12:44:55-0400
"E=\frac{V_2-V_1}{x}"

"E=\frac{23-21}{0.005 }=400\frac{V}{m}"